# The Role of Accuracy and Validation Effectiveness in Conversational Business Analytics

### v2

@@ Abstract
This study examines conversational business analytics, an approach that utilizes AI to address the technical competency gaps that hinder end users from effectively using traditional self-service analytics. By facilitating natural language interactions, conversational business analytics aims to empower end users to independently retrieve data and generate insights. The analysis focuses on Text-to-SQL as a representative technology for translating natural language requests into SQL statements. Developing theoretical models grounded in expected utility theory, the study identifies conditions under which conversational business analytics, through partial or full support, can outperform delegation to human experts. The results indicate that partial support, focusing solely on information generation by AI, is viable when the accuracy of AI-generated SQL queries leads to a profit that surpasses the performance of a human expert. In contrast, full support includes not only information generation but also validation through explanations provided by the AI, and requires sufficiently high validation effectiveness to be reliable. However, user-based validation presents challenges, such as misjudgment and rejection of valid SQL queries, which may limit the effectiveness of conversational business analytics. These challenges underscore the need for robust validation mechanisms, including improved user support, automated processes, and methods for assessing quality independently of end users’ technical competencies.
@@ 1 Introduction
Despite these benefits, self-service analytics has notable limitations. While end users are often domain experts, they often lack the technical skills required for advanced analytics tasks, such as navigating complex data structures, writing code, or using machine learning techniques. This skills gap can lead to errors and continued reliance on technical experts, undermining the autonomy and effectiveness of self-service analytics.
@@
Recent advances in generative AI, particularly the development of large language models, provide a transformative solution to these challenges. These models enable natural language interaction with analytics systems, removing the need for technical expertise. Building on this foundation, conversational business analytics is emerging as an innovative paradigm that redefines how users interact with business analytics systems. By leveraging large language models, conversational business analytics enables users to delegate tasks such as data retrieval, analysis and visualization to AI that understands and generates natural language output. This approach bridges the skills gap and extends access to sophisticated analytics tools to a wider range of organizational roles.
@@
There are a growing number of initiatives to extend self-service analytics with an AI-powered natural language interface. In response, software vendors are expanding their product portfolios to leverage individual data analytics. These solutions are designed to enable end users to seamlessly interact with data, generate reports and perform a wide range of analytical tasks independently.
@@
This study develops theoretical models, based on expected utility theory, to identify the conditions under which conversational business analytics using Text-to-SQL outperforms delegation to human experts. Central to this analysis is the interplay between accuracy (the ability of AI to generate correct information) and validation effectiveness (performance in correctly distinguishing between correct and incorrect information). The models examine two levels of AI support: partial support, where the AI generates information without additional validation, and full support, which includes a validation process to increase trustworthiness. The conditions under which conversational business analytics surpasses delegation to human experts are identified, with a particular focus on scenarios where validation should be performed to enhance trustworthiness and those where it should be omitted to avoid degradation in decision quality. These findings provide a structured framework for determining when and how AI-driven insight generation combined with validation can increase benefits while maintaining reliability, and offer practical guidance for implementing conversational business analytics.
@@
The structure of this study is as follows: Chapter 2 provides an overview of traditional self-service analytics, highlighting its limitations. Chapter 3 introduces conversational business analytics, exploring its transformative potential with a focus on large language models. This chapter also illustrates the challenges of information generation and subsequent validation, using Text-to-SQL as an example. Chapter 4 presents the models that explore the dynamics of partial and full support scenarios, focusing on their implications for information generation and validation. Finally, Chapter 5 concludes with a summary of the key findings, an acknowledgement of the limitations of the study, and a discussion of directions for future research.
@@ 2 Traditional Self-Service Analytics
Business analytics focuses on generating actionable insights to support data-driven decision making in business contexts [10, 18, 29, 80]. To achieve these goals, business analytics systems consist of components [31, 86, 65, 56] that automate four core tasks related to insight generation: data retrieval, data preparation, information generation and information visualisation (see Figure 1). In addition to these core tasks, there are complementary tasks such as ensuring data and information quality and establishing governance mechanisms. Although these complementary tasks are important for achieving the goals associated with business analytics, they are not discussed further in this study.
@@
Retrieving data involves collecting raw data from various source systems, such as structured data from relational databases, semi-structured data such as JSON or XML, and unstructured data such as text or images. The next step, preparing data, includes cleansing, transformation and enrichment processes to ensure quality and usability. Prepared data is stored in data warehouses and data marts for structured analysis, while semi-structured or unstructured data is stored in data lakes [45] or variants, prepared on demand.
@@
Generating information applies the three main methods of business analytics [17]: descriptive, predictive, and prescriptive analytics. Descriptive analytics summarizes historical and real-time data, providing insights into past and current performance using key performance indicators (KPIs), trend analysis, and target-performance comparisons [18]. Predictive analytics uses statistical models and machine learning techniques to forecast future developments, such as predicting customer churn or market trends. Prescriptive analytics recommends actions based on predictions, utilizing planning, simulation, and optimization methods to identify effective decisions. These methods are complementary, with prescriptive analytics relying on both descriptive and predictive results.
@@
Finally, visualizing information transforms analytical results into intuitive formats that support decision-making [71]. Effective visualizations, such as dashboards, simplify complex insights, enabling to explore trends, identify opportunities, and act quickly.
@@
Day-to-day information needs are typically met through predefined reports provided by business analytics systems. However, when new information needs arise, self-service analytics [30, 1, 47] enables end users to fulfill information needs independently, providing a viable alternative to the traditional reliance on expert (e.g. data engineers and data scientists) intervention. In this context, an end user refers to any organizational member who uses information to make decisions – whether to inform their own decisions or to prepare information for others, such as supervisors. These individuals are typically domain experts with extensive knowledge in their respective fields, distributed across various departments within the organization.
@@
The implementation of self-service analytics offers significant benefits [30, 1, 61]. It enables flexible and timely information generation, supporting early identification of opportunities and risks. This allows for faster decision making, both to capitalize on positive outcomes and to mitigate negative impacts. Self-service analytics also falls under the umbrella of ”’end-user computing” [59], where tasks traditionally managed by specialized departments are now performed by members of the organization. By eliminating the need to delegate information retrieval tasks to experts, self-service analytics removes dependencies and delays, thereby streamlining processes. In addition, it reduces the agency costs associated with delegation [27] by minimizing the exposure to hidden actions, which in turn reduces the need for incentives or oversight.
@@
Self-service analytics for descriptive purposes has traditionally relied on the data mart approach, enabling users to access predefined information and perform OLAP operations [14, 9] like drill-down and roll-up. However, data – customer, product, or process-related – exists often outside of data marts, in data warehouses, data lakes, or source systems. While real-time OLAP and in-memory computing have improved responsiveness, generating actionable insights often still requires expert assistance, leading to delays. To address this, greater flexibility is needed, allowing to independently retrieve and prepare data, perform analytics, and visualize results [1]. This autonomy enhances decision-making and supports agile responses to changing business needs.
@@
Effective use of self-service analytics requires technical expertise [68] that many end users lack. While end users typically have domain-specific knowledge, they often lack skills in critical areas such as data retrieval, modeling, machine learning, programming and navigating complex data structures [62]. For example, Microsoft’s financial data warehouse, comprising 632 tables, over 4,000 columns and 200 views [22], illustrates the significant challenges of navigating large and complex databases. This knowledge gap not only increases the likelihood of errors, but also limits the ability to independently generate actionable insights.
@@
Several studies underscore the importance of technical skills for the effective application of self-service analytics [30, 38, 5]. Imhoff and White highlight the necessity of user-friendly tools, noting that ”sophisticated analytics are often too intricate, complex, or difficult to construct for many information workers.” Similarly, Lennerholt, Laere, and Söderström emphasize the dual importance of intuitive tool design and comprehensive training programs to address the technical knowledge gap among non-technical users [38]. In an empirical study, Alparslan and Hügens examine the challenges faced by small and medium-sized enterprises in using business analytics systems [3]. Their findings show that for around 70% of respondents, inadequate technical skills are a key barrier to turning raw data into actionable insights.
@@ 3.1 Overview
Business analytics is undergoing significant development with the emergence of generative AI, giving rise to a new approach termed “Conversational Business Analytics” (CBA). This new paradigm leverages natural language processing to address persistent challenges in traditional self-service analytics, particularly the technical skill gap among end users. CBA enables natural language interactions for tasks such as data processing, analysis, and insight generation, presenting the potential to enhance the accessibility and efficiency of business analytics.
CBA shifts the focus from traditional graphical user interfaces to natural language-driven interactions, supported by advancements in large language models (LLMs) [79, 19, 7]. These models process natural language by tokenizing text and employing self-attention mechanisms to interpret contextual relationships. LLMs allow for context-aware and relevant outputs, with responses generated token by token based on probabilistic modeling. Trained on extensive text corpora, LLMs optimize billions of parameters to achieve high performance in natural language understanding and generation. As depicted in Figure 2, CBA facilitates the transformation of natural language inputs into actionable outputs, such as structured reports or visualizations.
CBA extends the capabilities of business analytics by automating core tasks for insight generation, including data retrieval, preparation, analytics, and visualization. It supports structured data (e.g., from data warehouses), semi-structured data (e.g., from data lakes), and unstructured data (e.g., textual documents). For structured and semi-structured data, semantic parsing techniques are used to convert natural language requests into executable code. For unstructured data, LLMs extract insights through advanced text processing methods. The Table 1 below summarizes core tasks and their corresponding references for recent advancements.
@@
Information Extraction and Generation [54], [72], [40]
@@
Table 1: Selected advancements in CBA regarding the core tasks of business analytics. These advancements are categorized based on their focus on specific core tasks of business analytics. It is important to note that many of these advancements affect multiple tasks simultaneously, highlighting the interconnected nature of tasks such as data retrieval, preparation, analysis, and insight generation.
CBA also introduces interactive exchanges between users and AI, enabling refinement of requests, validation of outputs, and clarification of insights. Such interaction underscores its user-centric and adaptive design, as exemplified by tools like OpenAI’s ChatGPT Advanced Data Analysis, which guide users through tasks such as analytical modeling and workflow optimization [72].
By integrating natural language processing, interactive communication, and support for diverse tasks and data types, CBA introduces a novel form of delegation in business analytics (see Figure 3). Through CBA, complex tasks such as data processing and information generation can be assigned to AI, enabling the autonomous generation and validation of insights on behalf of the end user. This delegation has the potential to reduce dependence on human experts while addressing the limitations associated with traditional self-service analytics.
Given the dynamic nature of CBA, this study focuses on Text-to-SQL as a key semantic parsing technology [34, 6, 4, 20, 89]. Text-to-SQL, a component of natural language interfaces to databases [42], translates natural language prompts into Structured Query Language (SQL) queries, enabling the retrieval and transformation of data from relational databases. This technology supports both data retrieval and the preparation of comprehensive workflows for generating actionable insights.
The origins of Text-to-SQL can be traced back to Codd’s pioneering work in the 1970s proposing natural language interfaces for ”casual users” to interact with relational databases [13]. Early implementations, such as rule-based systems [85], were limited in flexibility, but subsequent advancements, including Long Short-Term Memory (LSTM) networks [93], improved the handling of sequential inputs. More recently, transformer-based architectures have replaced LSTM networks, demonstrating superior performance for complex natural language processing tasks. Modern Text-to-SQL systems leverage LLMs for both training and inference, significantly improving accuracy and capability.
@@
The primary focus of this study is on the accuracy of Text-to-SQL, defined as its effectiveness in generating correct SQL queries. Ensuring accuracy is crucial, as errors in query generation can lead to incorrect insights and suboptimal decisions. Additionally, the study examines validation effectiveness, which reflects the ability to correctly distinguish between correct and incorrect information. Validation effectiveness plays a key role in verifying the reliability of generated information, particularly in high-stakes scenarios where flawed data can have significant consequences. The dual challenges of ensuring effectiveness in information generation and validation are explored in detail in the following sections.
@@ 3.2 Accuracy
Delegating tasks to human experts is grounded in their ability to correctly interpret and respond to information needs, as well as to build a shared understanding of the underlying goals and requirements [23, 32]. A similar dynamic exists when interacting with AI: the AI must understand the user’s request to generate the desired outcome. However, while natural language is flexible and often ambiguous, SQL is highly structured and formal. For instance, the key figure ”material availability” may be interpreted differently by the logistics and maintenance departments within the same organization due to the coexistence of multiple terminological systems [51, 87, 28]. Beyond terminological differences, ambiguities may also arise from the linguistic complexity of the request itself, such as vague formulations. To translate such requests into correct SQL queries, the AI must map the terms used in the request to the corresponding tables and columns within the data model (schema mapping). This requires a deep understanding of the semantics of the data model, including the specific meaning of fields and the relationships between tables. Thus, the AI must not only understand the request but also identify the relevant tables and columns that match the user’s inquiry.
The AI’s ability to interpret a natural language request and translate it into correct SQL code is referred to as its accuracy. Several distinct types of accuracy can be identified to assess different aspects of AI performance [34, 91, 76, 33]. Syntactic accuracy examines whether the SQL statements generated by the AI are executable. Execution accuracy evaluates whether the SQL query generated by the AI produces the expected result, even if the query’s structure differs from a reference SQL statement (often referred to as the ”gold standard”). In contrast, exact match accuracy is more stringent, as it requires not only the correct result but also that the generated SQL query exactly matches the gold standard.
Two main strategies have been proposed to enhance the accuracy of Text-to-SQL: fine-tuning and prompt design. Each of these methods addresses different aspects: fine-tuning focuses on adjusting the model parameters of the LLMs, while prompt design involves crafting the input to the LLM.
@@
    Fine-tuning: Training an LLM from scratch is resource-intensive, requiring vast amounts of data and computational power [66, 64, 74, 25]. As an alternative, organizations can leverage transfer learning, wherein a pre-trained LLM, which has already learned general language representations, is refined with domain-specific data. Fine-tuning involves adjusting the parameters of a pre-trained LLM to better align with specific organizational needs and domains. In the context of Text-to-SQL, fine-tuning involves training the model on labeled datasets consisting of pairs of natural language requests and corresponding SQL statements. This process helps the model understand the organization’s specific data models, schema, and terminology, enabling it to generate correct SQL queries based on user input.
    One major downside of fine-tuning is the potential for overfitting. This occurs when the model becomes overly specialized in domain-specific data and is less capable of adapting to other requests, making it harder to generalize to different contexts. Moreover, ongoing maintenance poses challenges: as data repositories such as data warehouses or marts evolve, schemas frequently change. Therefore, the fine-tuned LLM may require periodic updates to remain effective, increasing operational complexity and costs. Fine-tuning is not a one-time task; it requires continuous updates to avoid becoming outdated as the organizational data landscape evolves.
@@
    Prompt Design: Unlike fine-tuning, prompt design [25, 44, 64] does not alter the underlying parameters of the LLM. Instead, it leverages the model’s generalization capabilities by crafting prompts that guide the LLM toward correct SQL query generation. In few-shot learning [7, 44], for example, a user’s request is enriched with additional contextual information, such as data model details or examples of analogous SQL queries. The LLM processes the enriched prompt to infer the appropriate SQL query. Here, the LLM does not acquire new knowledge; rather, it uses its existing understanding along with the provided context to generate SQL queries that it has not previously encountered. Retrieval-Augmented Generation (RAG) further enhances prompt design by retrieving relevant contextual information, such as metadata stored in a vector database. Upon receiving user input, RAG integrates this context into the prompt, enabling the LLM to generate SQL queries. These queries undergo automated checks for syntax, schema alignment, and compliance with user requirements. If discrepancies are detected, the LLM’s self-correction mechanism refines the query. Once validated, the query is executed, and the information is presented to the user. Figure 4 illustrates the integration of LLMs with RAG, highlighting seamless context retrieval and query refinement.
@@
The integration of LLMs has significantly advanced the performance of Text-to-SQL. On the SPIDER benchmark [91], which assesses the generalization capabilities of Text-to-SQL models across a wide variety of databases, the performance of leading models has markedly improved, with execution accuracy increasing from approximately 54% to 91%. Exact match accuracy, which was around 5% initially, has risen to 82%. By comparison, the BIRD benchmark [41], which presents even greater complexity, reveals that current LLM-based Text-to-SQL models achieve execution accuracy of around 73%, underscoring the challenges posed by this benchmark.

### v3

@@ Abstract
This study examines conversational business analytics, an approach that utilizes AI to address the technical competency gaps that hinder end users from effectively using traditional self-service analytics. By facilitating natural language interactions, conversational business analytics aims to empower end users to independently retrieve data and generate insights. The analysis focuses on Text-to-SQL as a representative technology for translating natural language requests into SQL statements. Developing theoretical models grounded in expected utility theory, this study identifies the conditions under which conversational business analytics, through partial or full support, can outperform delegation to human experts. The results indicate that partial support, focusing solely on information generation by AI, is viable when the accuracy of AI-generated SQL queries leads to a profit that surpasses the performance of a human expert. In contrast, full support includes not only information generation but also validation through explanations provided by the AI, and requires sufficiently high validation effectiveness to be reliable. However, user-based validation presents challenges, such as misjudgment and rejection of valid SQL queries, which may limit the effectiveness of conversational business analytics. These challenges underscore the need for robust validation mechanisms, including improved user support, automated processes, and methods for assessing quality independent of the technical competency of end users.
@@ 1 Introduction
Despite these benefits, self-service analytics has several limitations. While end users are often domain experts, they often lack the technical skills required for advanced analytics tasks, such as navigating complex data structures, writing code, or using machine learning techniques. This skill gap can lead to errors and continued reliance on technical experts, thus undermining the autonomy and effectiveness of self-service analytics.
@@
Recent advances in generative AI, particularly the development of large language models, provide a transformative solution to these challenges. These models enable natural language interaction with business analytics systems, eliminating the need for technical expertise. Building on this foundation, conversational business analytics is emerging as an innovative paradigm that redefines how users interact with business analytics systems. By leveraging large language models, conversational business analytics enables users to delegate tasks such as data retrieval, analysis, and visualization to AI, which understands and generates natural language output. This approach bridges the skill gap and extends access to sophisticated analytics tools to a wider range of organizational roles.
@@
There is a growing number of initiatives to extend self-service analytics using AI-powered natural language interface. In response, software vendors are expanding their product portfolios to leverage individual data analytics. These solutions are designed to enable end users to interact seamlessly with data, generate reports and perform a wide range of analytical tasks independently.
@@
This study develops theoretical models, based on the expected utility theory, to identify the conditions under which conversational business analytics using Text-to-SQL outperforms delegation to human experts. Central to this analysis is the interplay between accuracy (the ability of AI to generate correct information) and validation effectiveness (the performance in correctly distinguishing between correct and incorrect information). The models examine two levels of AI support: partial support, in which the AI generates information without additional validation, and full support, which includes a validation process to increase trustworthiness. The conditions under which conversational business analytics surpass delegation to human experts are identified, with a particular focus on scenarios where validation should be performed to enhance trustworthiness and those where validation should be omitted to avoid degradation in information quality. These findings provide a structured framework for determining when and how AI-driven insight generation combined with validation can increase benefits while maintaining reliability, and offer practical guidance for implementing conversational business analytics.
@@
The structure of this study is as follows. Chapter 2 provides an overview of traditional self-service analytics, highlighting its limitations. Chapter 3 introduces conversational business analytics and explores its transformative potential with a focus on large language models. This section also illustrates the challenges of information generation and subsequent validation, using Text-to-SQL as an example. Chapter 4 presents models that explore the dynamics of partial and full support scenarios, focusing on their implications for information generation and validation. Finally, Chapter 5 concludes with a summary of the key findings, an acknowledgement of the limitations of the study, and a discussion of directions for future research.
@@ 2 Traditional Self-Service Analytics
Business analytics focuses on generating actionable insights to support data-driven decision-making in business [10, 18, 29, 80]. To achieve these goals, business analytics systems consist of components [31, 86, 65, 56] that automate four core tasks related to insight generation: data retrieval, data preparation, information generation and information visualization (see Fig. 1). In addition to these core tasks, there are complementary tasks such as ensuring data and information quality and establishing governance mechanisms. Although these complementary tasks are important for achieving goals associated with business analytics, they are not discussed further in this study.
@@
Retrieving data involves collecting raw data from various source systems, such as structured data from relational databases, semi-structured data such as JSON or XML, and unstructured data such as text or images. The next step, data preparation, includes cleansing, transformation and enrichment processes to ensure quality and usability. The prepared data are stored in data warehouses and data marts for structured analysis, whereas semi-structured or unstructured data are stored in data lakes [45] or variants, prepared on demand.
@@
Information generation applies the three main methods of business analytics [17]: descriptive, predictive, and prescriptive analytics. Descriptive analytics summarizes historical and real-time data, providing insights into past and current performance using key performance indicators (KPIs), trend analysis, and target-performance comparisons [18]. Predictive analytics uses statistical models and machine learning techniques to forecast future developments, such as predicting customer churn or market trends. Prescriptive analytics recommends actions based on predictions, utilizing planning, simulation, and optimization methods to identify effective decisions. These methods are complementary, with prescriptive analytics relying on both the descriptive and predictive results.
@@
Finally, visualizing information transforms analytical results into intuitive formats that support decision-making [71]. Effective visualizations, such as dashboards simplify complex insights, enable the exploration of trends, identify opportunities, and act quickly.
@@
Day-to-day information needs are typically met through predefined reports provided by business analytics systems. However, when new information needs arise, self-service analytics [30, 1, 47] enables end users to fulfill information needs independently, providing a viable alternative to the traditional reliance on expert (e.g. data engineers and data scientists) intervention. In this context, an end user refers to any organizational member who uses information to make decisions, whether to inform their own decisions or prepare information for others, such as supervisors. These individuals are typically domain experts with extensive knowledge in their respective fields, and are distributed across various departments within the organization.
@@
Self-service analytics offers significant benefits [30, 1, 61], enabling flexible and timely information generation to support the early identification of opportunities and risks. This allows for faster decision making, both to capitalize on positive outcomes and to mitigate negative impacts. Self-service analytics also falls under the umbrella of ”end-user computing” [59], where tasks traditionally managed by specialized departments are now performed by members of the organization. By eliminating the need to delegate information retrieval tasks to experts, self-service analytics removes dependencies and delays. In addition, it reduces agency costs associated with delegation [27] by minimizing exposure to hidden actions, which in turn reduces the need for incentives or oversight.
@@
Self-service analytics for descriptive purposes have traditionally relied on the data mart approach, enabling users to access predefined information and perform OLAP operations [14, 9], such as drill-down and roll-up. However, data – customer, product, or process-related – often exist outside of data marts, in data warehouses, data lakes, or source systems. Although real-time OLAP and in-memory computing have improved responsiveness, generating actionable insights often requires expert assistance, leading to delays. To address this, greater flexibility is needed, allowing the independent retrieval and preparation of data, performing analytics, and visualizing results [1]. This autonomy enhances decision-making and supports agile responses to changing business needs.
@@
The effective use of self-service analytics requires technical expertise [68], which many end users lack. While end users typically have domain-specific knowledge, they often lack skills in critical areas such as data retrieval, modeling, machine learning, programming and navigating complex data structures [62]. For example, Microsoft’s financial data warehouse, comprising 632 tables, over 4,000 columns and 200 views [22], illustrates the significant challenges of navigating large and complex databases. This knowledge gap not only increases the likelihood of errors, but also limits the ability to independently generate actionable insights.
@@
Several studies underscore the importance of technical skills in the effective application of self-service analytics [30, 38, 5]. Imhoff and White highlight the necessity of user-friendly tools, noting that ”sophisticated analytics are often too intricate, complex, or difficult to construct for many information workers.” Similarly, Lennerholt, Laere, and Söderström emphasize the dual importance of intuitive tool design and comprehensive training programs to address the technical knowledge gap among non-technical users [38]. In an empirical study, Alparslan and Hügens examine the challenges faced by small and medium-sized enterprises in using business analytics systems [3]. Their findings show that for around 70% of respondents, inadequate technical skills are a key barrier to turning raw data into actionable insights.
@@ 3.1 Overview
Business analytics is undergoing significant development with the emergence of generative AI, giving rise to a new approach termed “Conversational Business Analytics” (CBA). This new paradigm leverages natural language processing to address persistent challenges in traditional self-service analytics, particularly the technical skill gap among end users. CBA facilitates natural language interactions for core business analytics tasks, significantly improving their accessibility and effectivity.
CBA shifts the focus from traditional graphical user interfaces to natural language-driven interactions, supported by advancements in large language models (LLMs) [79, 19, 7]. These models process natural language by tokenizing text and employing self-attention mechanisms to interpret the contextual relationships. LLMs enable context-aware and relevant outputs by calculating the most probable response token by token. Trained on extensive text corpora, LLMs optimize billions of parameters to achieve high performance in natural language understanding and generation. As depicted in Fig. 2, CBA facilitates the transformation of natural language inputs into actionable outputs such as structured reports or visualizations.
CBA supports structured data (e.g., from data warehouses), semi-structured data (e.g., from data lakes), and unstructured data (e.g., textual documents). For structured and semi-structured data, semantic parsing techniques are used to convert natural language requests into executable code. For unstructured data, LLMs extract insights through advanced text processing methods. Table 1 summarizes the core tasks and their corresponding references for recent advancements.
Table 1: Selected advancements in CBA regarding the core tasks of business analytics.
@@
Information extraction & generation [54], [72], [40]
@@
CBA also introduces interactive exchanges between users and AI, enabling refinement of requests, validation of outputs, and clarification of insights. Such interaction underscores its user-centric and adaptive design, as exemplified by tools such as OpenAI’s ChatGPT Advanced Data Analysis, which guides users through tasks like training machine learning models and subsequent model optimization [72].
By integrating natural language processing, interactive communication, and support for diverse tasks and data types, CBA introduces a novel form of delegation in business analytics (see Fig. 3). Through CBA, complex tasks, such as data processing and information generation, can be assigned to AI, enabling autonomous generation and validation of insights on behalf of the end user. This delegation has the potential to reduce dependence on human experts, while addressing the limitations associated with traditional self-service analytics.
@@
Given the dynamic nature of CBA, this study focuses on Text-to-SQL [34, 6, 4, 20, 89] as a key semantic parsing technology. Text-to-SQL, a component of natural language interfaces to databases [42], translates natural language prompts into Structured Query Language (SQL) queries, enabling the retrieval and transformation of data from relational databases. This technology supports both data retrieval and the preparation of comprehensive workflows for generating actionable insights.
The origins of Text-to-SQL can be traced back to Codd’s pioneering work in the 1970s, proposing natural language interfaces for ”casual users” to interact with relational databases [13]. Early implementations, such as rule-based systems [85], were limited in flexibility, but subsequent advancements, including Long Short-Term Memory (LSTM) networks [93], have improved the handling of sequential inputs. More recently, transformer-based architectures have replaced LSTM networks, demonstrating superior performance in complex natural language processing tasks. Modern Text-to-SQL systems leverage LLMs for both training and inference, significantly improving accuracy and capability.
The primary focus of this study is on the accuracy of Text-to-SQL, which is defined as its effectiveness in generating correct SQL queries. Ensuring accuracy is crucial, because errors in query generation can lead to incorrect insights and suboptimal decisions. Additionally, this study examines validation effectiveness, which reflects the ability to correctly distinguish between correct and incorrect information. Validation effectiveness plays a key role in verifying the reliability of the generated information, particularly in high-stake scenarios where flawed data can have significant consequences. The dual challenges of ensuring the effectiveness of information generation and validation are explored in detail in the following sections.
@@ 3.2 Accuracy
Delegating tasks to human experts is grounded in their ability to correctly interpret and respond to information needs as well as to build a shared understanding of the underlying goals and requirements [23, 32]. A similar dynamic exists when interacting with AI. The AI must understand the user’s request to generate the desired outcome. However, while natural language is flexible and often ambiguous, SQL is highly structured and formal. For instance, the key figure ”material availability” may be interpreted differently by the logistics and maintenance departments within the same organization because of the coexistence of multiple terminological systems [51, 87, 28]. Beyond terminological differences, ambiguities may arise from the linguistic complexity of the request itself, such as vague formulations. To translate such requests into correct SQL queries, AI must map the terms used in the request to the corresponding tables and columns within the data model (schema mapping). This requires a deep understanding of the semantics of the data model, including the specific meaning of the fields and the relationships between tables. Thus, the AI must not only understand the request but also identify the relevant tables and columns that match the user’s inquiry.
The AI’s ability to interpret a natural language request and translate it into the correct SQL code is referred to as its accuracy. Various definitions of accuracy can be found in the literature, reflecting different perspectives and contexts [34, 91, 76, 33]. Syntactic accuracy examines whether the SQL statements generated by AI are executable. The execution accuracy evaluates whether the SQL query generated by the AI produces the expected result, even if the query’s structure differs from a reference SQL statement (often referred to as the ”gold standard”). In contrast, exact match accuracy is more stringent, as it requires not only the correct result but also that the generated SQL query exactly matches the gold standard.
Two main strategies have been proposed to enhance the accuracy of Text-to-SQL: fine-tuning and prompt design. Each of these methods addresses different aspects: fine-tuning focuses on adjusting the model parameters of a LLMs, whereas prompt design involves crafting the input to the LLM.
@@
    Fine-tuning: Training an LLM from scratch is resource-intensive and requires vast amounts of data and computational power [66, 64, 74, 25]. As an alternative, organizations can leverage transfer learning, wherein a pre-trained LLM that has already learned general language representations is refined with domain-specific data. Fine-tuning involves optimizing the parameters of a pre-trained LLM to enhance its performance on specific tasks and ensure better alignment with particular organizational needs and domain-specific requirements. In the context of Text-to-SQL, fine-tuning involves training the model on labeled datasets consisting of pairs of natural language requests and their corresponding SQL statements. This process helps the model to understand the organization’s specific data models, schema, and terminology, enabling it to generate correct SQL queries based on user input.
    A major drawback of fine-tuning is the risk of overfitting. This occurs when the model becomes overly specialized in domain-specific data and is less capable of adapting to other requests, making it more difficult to generalize to different contexts. Moreover, ongoing maintenance poses challenges: as data repositories (such as data warehouses or data marts) evolve, schemas frequently change. Therefore, the fine-tuned LLM may require periodic updates to remain effective, increasing operational complexity and costs. Fine-tuning is not a one-time task; it requires continuous updates to avoid becoming outdated as the organizational data landscape evolves.
@@
    Prompt Design: Unlike fine-tuning, prompt design [25, 44, 64] does not alter the underlying parameters of the LLM. Instead, it leverages the model’s generalization capabilities by crafting prompts that guide the LLM toward correct SQL query generation. In few-shot learning [7, 44], for example, a user’s request is enriched with additional contextual information, such as data model details or examples of analogous SQL queries. The LLM processes the enriched prompt to infer the appropriate SQL query. Here, the LLM does not acquire new knowledge; rather, it uses its existing understanding along with the provided context to generate SQL queries that it has not previously encountered. Retrieval-Augmented Generation (RAG) further enhances prompt design by retrieving relevant contextual information, such as metadata stored in a vector database. Upon receiving user input, RAG integrates this context into the prompt, enabling the LLM to generate SQL queries. These queries undergo automated checks for syntax, schema alignment, and compliance with user requirements. If discrepancies are detected, the LLM’s self-correction mechanism refines the query. Once validated, the query is executed, and the information is presented to the user. Fig. 4 illustrates the integration of LLMs with RAG, highlighting seamless context retrieval and query refinement.
@@
The integration of LLMs has significantly improved Text-to-SQL performance. On the SPIDER benchmark [91], which assesses the generalization capabilities of Text-to-SQL models across a wide variety of databases, the performance of leading models has markedly improved, with the execution accuracy increasing from approximately 54% to 91%. The exact match accuracy, which was initially approximately 5%, was increased to 82%. In comparison, the BIRD benchmark [41], which presents even greater complexity, reveals that the current LLM-based Text-to-SQL models achieve execution accuracy of around 73%, underscoring the challenges posed by this benchmark.